OpenStudy (anonymous):

HELP! calculate the lenght of the curve: x^(2/3)+y^(2/3)=9 thanks

OpenStudy (anonymous):

71 units...i guess..

OpenStudy (anonymous):

Well you need to get the limits first. Are they given? Get your equation into the form y=f(x) then apply the equation \[S= \int\limits_{a}^{b} \sqrt{(1+(dy/dx)^2) } dx\]

OpenStudy (kyosuke):

Mira, te recomiendo parametrizar tu ecuación de esta forma: x=27cos^3 (t) y=27sin^3(t) si los reemplazas uno con otro obtienes lo mismo ( x ^ (2/3) + y ^ (2/3) = 9) luego aplicas la integral para calcular la longitud \[4*\int\limits_{0}^{\pi/2} \sqrt {(27cos^{3} t)\prime^2+(27sen ^{3}t )\prime^2} dt\] si integras eso te debe salir 162 suerte.